OpenStudy (anonymous):

Multiply. (y-3)(y-5)

sam (.sam.):

y^2 - 30y +15

OpenStudy (anonymous):

\[y^2-8 y+15 \]

OpenStudy (anonymous):

(y-3)(y-5) = y^2 - 3y -5y + 15 = y^2 - 8y + 15